OpenStudy (anonymous):

Show that the series 1 - cos(1/n) converges absolutely ?

OpenStudy (anonymous):

let it be a function f(n)=1-cos(1/n) as 'n' increases and approach infinity, (1/∞) will become 0. cos(0)=1 therefore, 1-cos(1/∞)=1-cos(0)=1-1=0 the function converges to 0.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}(1-\cos (\frac{ 1 }{ n }))\]\[= (1-\cos (\frac{ 1 }{ \infty }))\]\[= (1-\cos (0))\]\[= (1-1)\]\[= 0\]

terenzreignz (terenzreignz):

@NicksonYap Yes, as a sequence, it is convergent, but what about as a series? :) \[\Large \sum_{n=1}^\infty \left[1-\cos\left(\frac1n\right)\right]\]

OpenStudy (anonymous):

ah, don't really know how to do this. The series is convergent because as 'n' increases, cos(1/n) gets closer and closer to 1, making 1-cos(1/n) closer and closer to 0. So the series will end at some point. Can we use integration instead?

terenzreignz (terenzreignz):

You can certainly try to integrate it. If you can prove that the integral is convergent, then the summation will also be convergent. LOL not all sequences that converge to zero are convergent as a series. Harmonic series, anyone? :) \[\Large \sum_{n=1}^\infty \frac1n \qquad \qquad \color{red}{\text{DIVERGENT!!!}}\]

OpenStudy (anonymous):

Why do we start from n=1? why not n=0?

terenzreignz (terenzreignz):

We simply can't start with n = 0? lol You try evaluating \[\Large 1 - \cos\left(\frac1n\right)\] when n = 0 I'll wait ;)

OpenStudy (anonymous):

You are right, the series can only start with n = 1 for its denominator is n. Anyway, is there anyway to change this series into a simple function which is possible to integrate?

terenzreignz (terenzreignz):

No idea. However...

terenzreignz (terenzreignz):

Are you panicking? :D

OpenStudy (anonymous):

My friend came up with the idea that \[1 - \cos \frac{ 1 }{ n } = \frac{ 1 - \cos(u) }{ 1 - u} \times (1-u) = (1 - u) sinu\]

OpenStudy (anonymous):

but i don't get it :))

terenzreignz (terenzreignz):

Because it doesn't make sense? lol

OpenStudy (anonymous):

it doesn't make sense to me because i don't know which equation is this. Do you know this or is it available ?

terenzreignz (terenzreignz):

Pay no heed to it. Use the Limit comparison test ^_^

OpenStudy (experimentx):

try expanding cos(1/n)

terenzreignz (terenzreignz):

@experimentX don't we have enough summations as it is? :D jk

OpenStudy (experimentx):

no no ... just approximate it. 1 - cos(1/n) < 1/n^2

terenzreignz (terenzreignz):

^ A bit hard to prove, methinks. LCT with 1/n^2 works ^_^

OpenStudy (anonymous):

but how can you come up with 1/n^2 ?

OpenStudy (experimentx):

you get that after expanding cos(1/n)

terenzreignz (terenzreignz):

oh... okay :(

terenzreignz (terenzreignz):

@sukanguyen To use the LCT, you have to have a suspicion of sorts, whether or not it's convergent, so suppose it's convergent.... We'll find a *known* convergent series with which to compare limits. We know the series \[\Large \sum_{n=1}^\infty \frac1{n^2}\] is convergent, so get this limit: \[\LARGE \lim_{x\rightarrow \infty}\frac{1 - \cos\left(\frac1x\right)}{\frac1{x^2}}\] If it goes to some number, then that will be enough to prove that your series is convergent ^_^

OpenStudy (experimentx):

well well ... that works too.

terenzreignz (terenzreignz):

Of course it does... I thought long and hard about it >:( ^_^

OpenStudy (anonymous):

if we can't start with n=0, why choose n=1? why not start with like n=0.00000001? It was not said that 'n' is discrete..

terenzreignz (terenzreignz):

It's a series, nick. A series is a sequence of partial sums. The domain of a sequence is the set of natural numbers. ^_^

OpenStudy (experimentx):

it's pointless to involve singularities due to zero.

terenzreignz (terenzreignz):

While we're at it @sukanguyen Try to evaluate that limit I gave you ^_^

OpenStudy (anonymous):

It's pretty hard :( as i prove,the limit approaches infinity

terenzreignz (terenzreignz):

No it isn't. In fact, it's indeterminate, heard of indeterminate forms? ^_^

OpenStudy (experimentx):

use L'hopital ... if you are doing sequence and series then I believe you have done L'hopital http://www.wolframalpha.com/input/?i=Limit%5Bx%5E2+%281++-+Cos%5B1%2Fx%5D%29%2C+x+-%3E+Infinity%5D although there are other ways you can do it.

OpenStudy (experimentx):

easy way is to expand cos(1/x) using Taylor series. the other way it to write 1 as cos(0) and use cos(C)-cos(D) formula .... you should end up with the same value.

OpenStudy (anonymous):

i know it's inderterminate. L'hospital is possible also

OpenStudy (anonymous):

Okay, i get it. Thank you guys :)